OpenStudy (anonymous):

pLEASE HELP factoring problem 4x-2(x-7)^2

OpenStudy (solomonzelman):

■ My interpretation of the problem: Factor, \(\color{#000000 }{ \displaystyle 4x-2(x-7)^2 }\) ■ You will need a couple rules to do this problem: \(\color{#000000 }{ \displaystyle (a-b)^2=a^2-2ab+b^2 }\) \(\color{#000000 }{ \displaystyle a(b+c+k)=a\cdot b+a\cdot c+a\cdot k=ab+ac+ak }\)

OpenStudy (solomonzelman):

And then, after applying these two rules to your problem (that is - after expanding), you need to add like terms...

OpenStudy (anonymous):

do you still not get it?